OpenStudy (anonymous):

[9.06] Rebekah kicks a soccer ball off the ground and in the air with an initial velocity of 25 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 9.5 feet 9.8 feet 10.2 feet 10.7 feet

OpenStudy (anonymous):

Need help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

You are basically trying to find the x value of the turning point of the parabola.

OpenStudy (anonymous):

I meant y sorry!

OpenStudy (anonymous):

will it be B then i don't know

OpenStudy (anonymous):

H(t)= -16t^2+25t+0 So the maximum height is at the vertex. t = -b/2a= -25/ (2*-16)= 25/32 sec or about .78sec Plug into H(t) H(25/32)= -16(.78)^2+25(.78)= 9.77 ft or about 9.8 ft

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

Yeah you got it :)

OpenStudy (anonymous):

cool thanks

OpenStudy (anonymous):

can you help on this... [9.01] Determine whether the graph of y = x2 − 4x + 1 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (2, -3) Maximum; (2, -3) Minimum; (-3, 2) Maximum; (-3, 2)

OpenStudy (anonymous):

The x^2 is positive so...?

OpenStudy (anonymous):

its up

OpenStudy (anonymous):

Yeah and that means?

OpenStudy (anonymous):

its Minimum

OpenStudy (anonymous):

That's it

OpenStudy (anonymous):

What would you do next?

OpenStudy (anonymous):

Now its either a or c.

OpenStudy (anonymous):

i thought it be A

OpenStudy (anonymous):

That's correct.

OpenStudy (anonymous):

You already did the -b/2a and substituted the result in the equation.